# Revisiting Global Income Convergence in the 21st Century

### v1

@@ Appendix B Robustness and Additional Exercises
B.1 Unconditional Convergence with 1960 Base Year
Table 5: Beta Convergence for 1980-2000 and 2000-2019.
All Countries Outside SSA
(1) (2) (3) (4)
1980-2000 2000-2019 1980-2000 2000-2019
$\beta$ 0.002 --0.009∗∗∗ --0.004 --0.014∗∗∗
(0.003) (0.003) (0.002) (0.004)
N 107 107 72 72
R2 0.007 0.090 0.034 0.157
Standard errors in parentheses
* $p<0.10$, ** $p<0.05$, *** $p<0.01$
Table 5 presents the results of $\beta$-convergence regressions, using 1960 as the base year for initial income levels. The key objective here is to determine whether the choice of base year affects the conclusions regarding income convergence.
The results confirm that the findings are robust to the use of a different base year. For both the full sample of countries and the sub-sample excluding Sub-Saharan Africa (SSA), there is no evidence of convergence between 1980 and 2000. In this period, the $\beta$ coefficient is close to zero and statistically insignificant in both samples, indicating no relationship between initial income levels and subsequent growth rates.
In contrast, the period 2000-2019 shows strong evidence of convergence, with statistically significant negative $\beta$ coefficients for both the full sample and the sub-sample excluding SSA. The magnitude of convergence is similar to that found when using 1980 as the base year. For the full sample of countries, the $\beta$ coefficient is --0.009, while for countries outside SSA, it is --0.014. These results suggest that, post-2000, poorer countries have been catching up with richer countries at similar rates, regardless of whether 1960 or 1980 is used as the starting point for income levels.
B.2 Robustness to different measures of GDP
Table 6: Beta Convergence for 2000-2019 (Excluding SSA)
RGDPO RGDPE CGDPO CGDPE
$\beta$-0.015∗∗∗ --0.014∗∗∗ --0.015∗∗∗ --0.014∗∗∗
(0.003) (0.003) (0.003) (0.003)
N 91 91 91 91
R2 0.238 0.221 0.250 0.221
Note: p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01
The Penn World Table (PWT) provides multiple measures of GDP, converted into common currency units, using different approaches to national income accounting. Two measures of real GDP are reported: one based on the output side (RGDPO) and another based on the expenditure side (RGDPE). Similarly, two measures of current-price GDP are available: the output-side measure (CGDPO) and the expenditure-side measure (CGDPE).
Table 6 presents a robustness check of the $\beta$-convergence results for the period 2000-2019, using different measures of per capita GDP for countries outside Sub-Saharan Africa. Each column corresponds to a different measure of GDP used in the calculation of per capita income.
The results indicate that the rate of convergence remains remarkably stable across all four GDP measures. In each case, the $\beta$ coefficient is negative and statistically significant at the 1% level, with convergence rates ranging between --0.014 and --0.015. This consistency suggests that the pattern of income convergence for countries outside Sub-Saharan Africa is robust to the choice of GDP measure, whether based on real output, real expenditure, or current-price accounting. The $R^{2}$ values also indicate a similar explanatory power across all specifications, with values ranging from 0.221 to 0.250.
B.3 Income difference for all countries
Table 7: Income Differences across Years (Full Sample)
Year P90/P10 P90/P50 P50/P10 Var(log(GDPpc)) Income Ratio
1980 18.56 5.30 3.50 1.15 18.49
1990 24.95 5.55 4.50 1.37 23.26
2000 31.95 6.88 4.64 1.62 29.53
2010 27.63 4.51 6.12 1.54 30.71
2019 25.06 4.41 5.69 1.47 27.17
@@ B.3 Income difference for all countries
Table 7 presents income differences across various measures for the full sample of countries, complementing the analysis of countries outside Sub-Saharan Africa (SSA) provided in the main text. The table reports several key indicators of income dispersion, including the P90/P10, P90/P50, and P50/P10 percentile ratios of GDP per capita, the variance of the log of GDP per capita, and the income ratio between the top 5 and bottom 5 of countries.
A key observation from this table is that the decline in income disparities post-2000 has been relatively modest for the full sample of countries, particularly compared to the more substantial reductions observed in countries outside SSA. For instance, the P90/P10 ratio fell from 31.95 in 2000 to 25.06 in 2019, indicating some reduction in inequality between the richest and poorest economies. However, this decline is less pronounced compared to the convergence seen outside SSA, as shown in the main text.
Notably, the P50/P10 ratio increased from 4.64 in 2000 to 5.69 in 2019, reflecting a widening income gap between middle-income and poor economies. This increase is largely driven by the stagnation of many Sub-Saharan economies, which have struggled to keep pace with global growth. Consequently, while income disparities have narrowed somewhat at the top end of the distribution (as seen in the decline in the P90/P10 and P90/P50 ratios), the lower half of the income distribution has experienced greater divergence, especially between middle-income and poorer countries.
B.4 Accounting for Income Convergence (Full Sample)
Table 8: Accounting for Income Convergence (Full Sample)
Year $\Delta_{t}\log(y(p_{n})/y(p_{0}))$ Contribution Contribution Contribution
(%) A (%) K/Y (%) $h$ (%)
Panel A: 1980-2000
P90/P10 54 131 -24 -6
P90/P50 15 247 -67 -87
P50/P10 39 85 -10 23
Panel B: 2000-2019
P90/P10 -27 81 0 19
P90/P50 -29 48 24 24
P50/P10 3 -233 267 67
Panel C: 1980-2019
P90/P10 27 181 -48 -30
P90/P50 -14 -164 121 143
P50/P10 42 62 10 26
Table 8 presents the accounting decomposition for income differences across the global income distribution for three periods: 1980-2000, 2000-2019, and 1980-2019. The analysis includes all economies, including Sub-Saharan Africa.
Panel A of Table 8 shows that between 1980 and 2000, income disparities widened across the income distribution. The P90/P10 income ratio increased by 54%, indicating divergence between the richest and poorest countries. TFP differences contributed significantly to this divergence, accounting for 131% of the increase. Convergence in the capital-output ratio and human capital partially offset this effect, contributing -24% and -6%, respectively. Similarly, the P90/P50 ratio increased by 15%, with TFP differences contributing 247% to divergence, while capital-output ratios and human capital contributed negatively, offsetting some of the divergence. The P50/P10 ratio increased by 39%, mainly due to TFP differences (85%), with minimal offset from inputs.
Panel B of Table 8 reveals that between 2000 and 2019, income disparities decreased between the richest and poorest countries. The P90/P10 income ratio declined by 27%, indicating convergence. TFP differences contributed 81% to this convergence, while human capital contributed 19%. The capital-output ratio had a negligible effect in this case. The P90/P50 ratio decreased by 29%, with TFP and inputs contributing equally to convergence. Notably, the P50/P10 ratio showed a slight divergence of 3%. This divergence was due to a negative contribution from TFP (-233%), but substantial convergence in the capital-output ratio (267%) and human capital (67%) more than compensated for this, indicating significant input-driven convergence among middle-income and poor countries.
Panel C of Table 8 presents the findings for the entire period from 1980 to 2019. Over this timeframe, the P90/P10 income ratio increased by 27%, indicating overall divergence between the richest and poorest countries. TFP differences were the main driver, contributing 181% to this divergence. Convergence in the capital-output ratio (-48%) and human capital (-30%) partially mitigated the divergence caused by TFP. The P90/P50 ratio decreased by 14%, reflecting convergence between the richest and middle-income countries. This convergence was driven by improvements in the capital-output ratio (121%) and human capital (143%), which offset the divergence from TFP differences (-164%). The P50/P10 ratio increased by 42%, indicating divergence between middle-income and poor countries, primarily due to TFP differences (62%).
B.5 Variance Decomposition
@@
This section complements the main accounting analysis by performing the accounting based on variance as a measure of income dispersion, which is more commonly used in the literature. However, I prefer the percentile decomposition over variance decomposition for several reasons. First, percentile ratios allow for continuous variation in parameters across time and countries. Second, percentile ratios are less sensitive to outliers compared to variance measures. Furthermore, variance decomposition includes a term for the covariance between total factor productivity (TFP) and inputs, which can be difficult to interpret. For this exercise, I use a Cobb-Douglas production function, and a capital share parameter ($\alpha$) of 0.46, which is the observed mean in PWT10.01:
@@
$\frac{Y}{L}=A(\frac{K}{Y})^{\frac{\alpha}{1-\alpha}}(h)$
$y=Ay^{kh}$
where $y=\frac{Y}{L}$, $y^{kh}=\frac{K}{Y})^{\frac{\alpha}{1-\alpha}}(h)$. Consequently,
$Var(\ln(y))=Var(\ln(A))+Var(\ln(y^{kh}))+2Cov(\ln(A),\ln(y^{kh}))$
Table 9: Accounting for Variance Decline
Years $\Delta_{t}Var(\ln(y))$ $\Delta_{t}Var(\ln(A))$ $\Delta_{t}Var(\ln(y^{kh}))$ $\Delta_{t}$ 2$\times$×Cov(A,$y^{kh}$)
Outside SSA
1980-2000 11 141 -126 86
2000-2019 -39 27 36 36
1980-2019 -27 -20 104 16
Full Sample
1980-2000 51 69 -20 51
2000-2019 -19 66 43 -9
1980-2019 31 71 -59 88
Table 9 presents the variance decomposition of log income per worker ($\ln(y)$) for both countries outside Sub-Saharan Africa (SSA) and the full sample over three periods: 1980-2000, 2000-2019, and 1980-2019.
For countries outside SSA, between 1980 and 2000, the variance of log income per worker increased by 11%, indicating a slight divergence in income levels among these countries. This increase was primarily driven by a substantial rise in the variance of log TFP, which surged by 141%, suggesting that differences in total factor productivity widened significantly during this period. The covariance term also increased by 86%, further contributing to divergence. However, the variance of log inputs decreased by 126%, indicating strong convergence in physical and human capital accumulation. Had it not been for the convergence in inputs, income variance would have been much higher, as the divergence from TFP differences and the covariance effect would have led to an even greater increase in income disparities.
Between 2000 and 2019, the variance of log income per worker outside SSA decreased by 39%, reflecting a significant convergence in income levels. Decline in variance of inputs accounted for 36% of convergence. The contribution of inputs increases to 54% when including half the covariance term.
The role of inputs is even more pronounced for the time period 1980-2019. During this period the variance of income decline by 27%, with decline in variance of inputs accounting for 104% of the decline, while TFP diverged during this period.
This shows that inputs account for almost half of decline since 2000, and almost all of decline since 1980 in variance of income outside SSA.
For the full sample of countries, between 1980 and 2000, the variance of log income per worker increased by 51%. This increase was primarily driven by a rise in the variance of log TFP by 69%. The covariance term also increased by 51%, further contributing to divergence. However, the variance of log inputs decreased by 20%. Between 2000 and 2019, the variance of log income per worker in the full sample decreased by 19%. Reduction in variance of TFP contributed to 66% of the decline, while reduction in variance of inputs contributed to 43% of decline. Covariance term increased by 9%, contributing to divergence. Over the entire period from 1980 to 2019, the variance of log income per worker in the full sample increased by 31%, indicating a net divergence in global income levels over the long term. This divergence was primarily driven by a substantial rise in the variance of log TFP, which increased by 71%, and an increase in the covariance term by 88%, both significantly contributing to divergence. However, the variance of log inputs decreased by 59%, indicating strong convergence in physical and human capital accumulation.
B.6 Capital Accumulation in Sub-Saharan Africa
It is puzzling to observe a persistent lack of convergence in Sub-Saharan African economies. Despite substantial improvements in policy outcomes across the region, as documented by [6], this has not translated into faster economic growth compared to economies outside Sub-Saharan Africa. [6] measures poor policy outcomes using indicators such as high inflation, significant black market premiums, and negative interest rates, among others. While these measures have improved considerably, growth in Sub-Saharan Africa remains sluggish relative to other regions. Although a comprehensive investigation into the reasons behind this lack of convergence is beyond the scope of the current study, I find that the slow rate of input accumulation is likely a key contributing factor. [17] further highlight capital flight from Sub-Saharan economies to the rest of the world, driven by economic and political uncertainties, which may exacerbate the region’s stagnation.
Figure 2: Capital-Output Ratio across regions
Figure 2 illustrates the mean capital-output ratio, weighted by population size, across different regions. While the capital-output ratio has continued to rise in regions outside Sub-Saharan Africa, Sub-Saharan Africa is the only region where a decline in the capital-output ratio is observed after 2000. This decline likely reflects the region’s challenges in accumulating capital, which may be linked to political instability, economic uncertainty, and capital flight. The capital-output ratio in Figure 2 is computed using a sample of countries for which investment data has been available since 1970 or earlier, ensuring consistent comparisons across regions.
B.7 Influence of Venezuela
Figure 3: Log GDP per capita (2017\$ PPP) for Venezuala
Figure 3 illustrates the significant decline in Venezuela’s GDP per capita, measured in 2017 PPP dollars, following 2010. The sharp drop in Venezuela’s GDP per capita is a direct consequence of the country’s political and economic crisis. Given the unique circumstances surrounding Venezuela’s economic decline, it stands as an outlier in cross-country analyses. Therefore, I exclude Venezuela from the calculation of income variance, as variance is particularly sensitive to outliers.

### v2

@@ Appendix B Robustness and Additional Exercises
B.1 Robustness to different measures of GDP
Table 3: Beta–Convergence Coefficients Across GDP Measures
GDP Measure
RGDPO RGDPE
Panel A: All Countries
Current sample $\beta$ –0.0062 –0.0052
(SE) (0.0017) (0.0016)
$N$ 155 155
PSS sample $\beta$ –0.0052 –0.0043
(SE) (0.0016) (0.0016)
$N$ 124 124
Panel B: Outside SSA
Current sample $\beta$ –0.0150 –0.0128
(SE) (0.0023) (0.0022)
$N$ 115 115
PSS sample $\beta$ –0.0155 –0.0135
(SE) (0.0021) (0.0021)
$N$ 90 90
Notes: $\beta$ coefficients are estimated by nonlinear least squares as in (1), with robust standard errors in parentheses. RGDPO and RGDPE denote output-side and expenditure-side real GDP per capita at chained PPPs (million 2017 US\$) from PWT 10.01. PSS sample refers to the sample of country used in [17].
@@ B.1 Robustness to different measures of GDP
Across all countries the estimated convergence rate lies between $-0.0062$ (output-side GDP, current sample) and $-0.0043$ (expenditure-side GDP, PSS sample); all coefficients are negative and statistically significant at the one-percent level. Restricting attention to economies outside SSA roughly triples the speed of convergence, with point estimates ranging from $-0.0155$ to $-0.0128$. It can thus be concluded that $\beta$-income convergence remains robust to the choice between RGDPO and RGDPE.
Appendix C Additional Figures
Figure 6: Decomposition Analysis (excluding SSA): P90/P50
Notes: The figure decomposes the change in the log P90/P50 output ratio into the contributions of total factor productivity, the capital-output ratio, and human capital for the periods 1980-2000 and 2000-2019 for countries outside Sub-Saharan Africa. Inputs (capital-output ratio and human capital) account for 66% of the convergence in the log P90/P50 ratio between 2000-2019 and for the entire convergence over 1980-2019.
Figure 7: Decomposition Analysis (excluding SSA): P50/P10
Notes: The figure decomposes the change in the log P50/P10 output ratio into the contributions of total factor productivity, the capital-output ratio, and human capital for the periods 1980-2000 and 2000-2019 for countries outside Sub-Saharan Africa. Inputs (capital-output ratio and human capital) account for 32% of the convergence in the log P50/P10 ratio between 2000-2019 and for 70% of the convergence over 1980-2019.
Figure 8: Decomposition Analysis (full sample): P90/P50
Notes: The figure decomposes the change in the log P90/P50 output ratio into the contributions of total factor productivity, the capital-output ratio, and human capital for the periods 1980-2000 and 2000-2019 using the full sample of countries. Inputs (capital-output ratio and human capital) account for 49% of the convergence in the log P90/P50 ratio between 2000-2019 and for the entire convergence over 1980-2019.
Figure 9: Decomposition Analysis (full sample): P50/P10
Notes: The figure decomposes the change in the log P50/P10 output ratio into the contributions of total factor productivity, the capital-output ratio, and human capital for the periods 1980-2000 and 2000-2019 using the full sample of countries. During 2000-2019 inputs (capital-output ratio and human capital) diverged, whereas TFP exhibited a slight convergence, yielding a net divergence in the P50/P10 ratio. Over the entire horizon 1980-2019 TFP divergence explains 63 % of the overall income divergence, while input divergence accounts for a further 27 %.
Figure 10: Decomposition Analysis (excluding SSA, $\alpha=1/3$): P90/P50
Notes: The figure decomposes the change in the log P90/P50 output ratio into the contributions of total factor productivity, the capital-output ratio, and human capital for 1980–2000 and 2000–2019, assuming a constant capital-income share of one-third. Inputs (capital–output ratio and human capital) account for 47% of the convergence in 2000–2019 and for the entire convergence over 1980–2000.
Figure 11: Decomposition Analysis (excluding SSA, $\alpha=1/3$): P50/P10
Notes: The figure decomposes the change in the log P50/P10 output ratio into the contributions of total factor productivity, the capital–output ratio, and human capital for 1980–2000 and 2000–2019, assuming a constant capital-income share of one-third. Inputs account for 10% of the convergence in 2000–2019 and for 37% of the convergence over 1980–2000.
Figure 12: Decomposition Analysis (full sample, $\alpha=1/3$): P90/P10
Notes: The figure decomposes the change in the log P90/P10 output ratio into the contributions of total factor productivity, the capital–output ratio, and human capital for 1980–2000 and 2000–2019, assuming a constant capital-income share of one-third. Capital convergence accounts for 28% of the convergence in 2000–2019. Over the full 1980–2019 horizon the ratio diverges overall, driven by large TFP divergence in 1980–2000; capital convergence mitigates but does not overturn this early divergence.
Figure 13: Decomposition Analysis (full sample, $\alpha=1/3$): P90/P50
Notes: The figure decomposes the change in the log P90/P50 output ratio into the contributions of total factor productivity, the capital–output ratio, and human capital for 1980–2000 and 2000–2019, assuming a constant capital-income share of one-third. Inputs account for 42.6% of the convergence in 2000–2019 and for the entire convergence over 1980–2019.
Figure 14: Decomposition Analysis (full sample, $\alpha=1/3$): P50/P10
Notes: The figure decomposes the change in the log P50/P10 output ratio into the contributions of total factor productivity, the capital–output ratio, and human capital for 1980–2000 and 2000–2019, assuming a constant capital-income share of one-third. In 2000–2019 overall output diverges despite TFP convergence because of divergence in capital inputs. Over 1980–2019 both TFP and capital diverge, reinforcing the long-run divergence in the P50/P10 ratio.
C.1 Variance Decomposition
@@
This section complements the main accounting analysis by performing the accounting based on variance as a measure of income dispersion, which is more commonly used in the literature. However, I prefer the percentile decomposition over variance decomposition for several reasons. First, percentile ratios allow for continuous variation in parameters across time and countries. Second, percentile ratios are less sensitive to outliers compared to variance measures. Furthermore, variance decomposition includes a term for the covariance between total factor productivity (TFP) and inputs, which can be difficult to interpret. For this exercise, I use a Cobb-Douglas production function, and a capital share parameter ($\alpha$) of 0.46, which is the observed mean in PWT10.01.
@@
$\frac{Y}{L}=A\left(\frac{K}{Y}\right)^{\frac{\alpha}{1-\alpha}}h,$
or
$y=A\,y^{kh},$
where $y=\frac{Y}{L}$ and $y^{kh}=\left(\frac{K}{Y}\right)^{\frac{\alpha}{1-\alpha}}h.$ Consequently,
$Var(\ln(y))=Var(\ln(A))+Var(\ln(y^{kh}))+2\,Cov(\ln(A),\ln(y^{kh})).$
Figure 15: Variance Decomposition Outside Sub-Saharan Africa
Notes: The figure decomposes the change in the variance of log output per worker into the contributions of total factor productivity (TFP), the combined variance of physical- and human-capital inputs, and the covariance between TFP and inputs for the intervals 1980-2000 and 2000-2019 for countries outside Sub-Saharan Africa.
Figure 16: Variance Decomposition, Full Country Sample
Notes: The figure reports the same variance decomposition for the full set of countries.
Figures 15 and 16 trace how the dispersion of income per worker has evolved by separating the variance of log output into three components: the variance of total factor productivity, the variance of combined physical- and human-capital inputs, and their covariance.
For countries outside Sub-Saharan Africa, the variance of log output per worker rose by 11 percent between 1980 and 2000. TFP divergence contributed 141 percent of this increase, and the covariance between TFP and inputs added a further 86 percent, whereas convergence in physical and human capital offset 126 percent of the potential rise. In the subsequent period, 2000-2019, the variance fell by 39 percent: convergence in inputs accounted for 36 percent of that reduction, the fall in the covariance term accounted for another 36 percent, and TFP convergence accounted for the remainder. Taken together, these forces produced a net 27-percent decline in the variance over 1980-2019, with input convergence contributing 104 percent of the total reduction.
For the full global sample, the variance of log output increased by 51 percent between 1980 and 2000, with TFP divergence responsible for 69 percent of the rise and the covariance term for 51 percent, while input convergence offset 20 percent. During 2000-2019 the variance contracted by 19 percent: TFP convergence contributed 66 percent of the decline, input convergence 43 percent, and a further expansion of the covariance term offset 9 percent. Over the entire horizon 1980-2019 the variance remained 31 percent above its initial level; TFP divergence accounted for 71 percent of this long-run increase and the covariance term for 88 percent, whereas input convergence reduced the total by 59 percent.
Figure 17: Log GDP per capita (2017\$ PPP) for Venezuela
Notes: The series displays a pronounced decline in Venezuela’s GDP per capita after 2010, reflecting the country’s political and economic crisis; Venezuela is treated as an outlier and excluded from the income-variance calculations.